Find the present value of the following annuities. Assume the discounting occurs once a year A. The customer agreed to pay rent at the end of each year and the landlord charges $21,600 per year . The required rate of return of the landlord is 6.85%. The rent contract will last for 7 years Required: Find the present value of this agreement for the landlord B. A customer approached a car sales agent and makes a car lease agreement for 6 years. The cash flow [annual] is $6000. The agent charges 7.2% interest on such contract Required: How much will be the present value of such an agreement ?


1
Expert's answer
2021-05-27T13:48:34-0400

(a)Present value of annuity"=Annual\space Cash\space flow×(\frac{1−(1+r)^{−n}}{r})"

r is rate of return

n is time period

"=\$21,600\times (\frac{1−(1+0.0685)^{−7}}{0.0685})"

"=\$117,019.89"

(B)

Present value of annuity"=Annual\space Cash\space flow×(\frac{1−(1+r)^{−n}}{r})"

r is rate of return

n is time period

"=\$6000\times (\frac{1−(1+0.072)^{−6}}{0.072})"

"=\$28,423.51"